If 55.60mL of 0.2221M HCl was needed to titrate 30.00 of NaOH to its equivalence point. What is the molar concentration of NaOH
1
Expert's answer
2020-10-10T08:35:42-0400

Balanced reaction between HCl and NaOH is

HCl + NaOH --------> NaCl + H2O

nA = mole of HCl =1

nB = mole of NaOH =1

(MAVA/nA) = (MBVB/nA)

A represent HCl

B represent NaOH

{(0.2221M)×(55.60mL)/1} = (MB)(30)/1

MB = {(0.2221M)(55.60)/30mL}

MB = 0.411M NaOH

MB = molar Concentration of NaoH

MB = 0.411 M NaOH